# Minimmit: Fast Finality with Even Faster Blocks

### v6

@@ 6.3. Erasure coding
Implementation for Minimmit:
    Set $d=f+1$ (ensuring availability despite $f$ Byzantine failures).
    Leader sends one unique fragment to each processor.
    Communication overhead: $n(\mathtt{block\_size}/d)=n(\mathtt{block\_size}/(f+1))\approxeq 5\mathtt{block\_size}$.
    Processors vote only after receiving and validating their fragment.
    Each vote includes the processor’s fragment to enable block reconstruction.
Correctness guarantee. Since any notarization requires votes from at least $2f+1$ processors, and at most $f$ are Byzantine, at least $f+1$ correct processors must contribute to each notarisation, meaning that they send their fragment to all processors. This ensures sufficient fragments are available for all processors to reconstruct any notarised block.
@@ 6.4. Threshold signatures and communication complexity
Section 4 assumes correct processors send newly received transactions to all others. In practice, transactions are typically disseminated through gossip networks, where each processor forwards transactions to a small, constant number of peers. If transactions are constant-bounded in size, this approach maintains constant communication overhead per processor per transaction. Alternative dissemination mechanisms like Narwhal (Danezis et al., 2022) can also be employed. In our analysis here, we treat transaction propagation as a black box and focus on the consensus layer. We assume blocks have constant-bounded size and, following standard practice, analyse only the communication costs required for consensus, taking mempool formation as given.
@@
Communication complexity. Each notarisation and nullification contains $\Omega(n)$ signatures, resulting in communication complexity greater than $O(n^{2})$ per view. The standard approach of using threshold signatures (Boneh et al., 2001; Shoup, 2000) can be used to ensure that the communication complexity per view is $O(n^{2})$. We redefine:
  • An ‘M-notarisation’ for $b$ to be threshold signature (of constant-bounded length for a given security parameter) formed from $2f+1$ votes for $b$ by different processors.
  • A ‘nullification’ for view $v$, to be a threshold signature formed from $2f+1$ nullify(v)(v) messages by different processors.
Threshold signature implemetation details. If $p_{i}$ has not already formed or received an M-notarisation for $b$, then, upon receipt of $2f+1$ votes for $b$ by different processors, $p_{i}$ combines the $2f+1$ signatures to form an M-notarisation (a single threshold signature). Rather than storing and forwarding the $2f+1$ votes, $p_{i}$ stores and forwards the M-notarisation. Similarly, $p_{i}$ stores and sends nullifications, rather than storing and sending large collections of nullify(v)(v) messages. Since partial signatures for any processor can be derived from $2f+1$ partial signatures, we must now stipulate that processors finalise a block $b$ upon receiving nfn-f votes for $b$ directly from the corresponding processors.
@@ 7. Experiments
Latency and bandwidth model. For every message transmission we sample a delay from a normal distribution with mean equal to the p50171717https://www.cloudping.co/api/latencies?percentile=p_50&timeframe=1Y AWS inter-region latency and standard deviation $(\text{p90}-\text{p50})$. [^18] https://www.cloudping.co/api/latencies?percentile=p_90&timeframe=1Y Each processor has symmetric 1 Gbps (125,000,000 B/s) ingress and egress budgets, so message delivery time is the sum of the sampled network delay and the bandwidth-limited transmission time. We assume linear bandwidth usage and max-min fairness for bandwidth allocation, yielding conservative latency estimates. In practice, burstable bandwidth and traffic prioritisation provided by cloud operators would only decrease the reported latencies.
@@
Introducing a naïve Reed–Solomon erasure coding scheme reduces this bottleneck. We split each block into 50 fragments, have the leader broadcast a fragment to every processor, and require replicas to broadcast their fragment to all other processors when casting a vote. The fragment size is determined by the quorum needed for view progression: Simplex targets $3f+1$ replicas and therefore transmits 61.72 KB fragments so that any $f+1$ fragments suffice for reconstruction; Minimmit targets $5f+1$ replicas and sends 104.86 KB fragments to ensure $f+1$ fragments suffice; Kudzu also targets $5f+1$ replicas but requires 55.19 KB fragments to guarantee reconstruction from $2f+1$ fragments. Table 6 reports the resulting latencies, increasing Minimmit’s effective processing rate from 10,000 to 25,000 transactions per second at a lower transaction latency.
@@
Minimmit $216.0\pm 21.46$ ms $290.32\pm 28.52$ ms $506.28\pm 7.07$ ms
With coding enabled, Minimmit achieves the lowest view latency but not the lowest transaction latency, despite each processor only emitting a single vote message per view (rather than driving concurrent paths). This result is a product of the smaller fragment size required by Kudzu (47.37% smaller than Minimmit’s), which offsets Minimmit’s lower threshold for view progression. With the lowest view latency, Minimmit still maximises throughput across all protocols, since blocks of a fixed size are produced at a faster rate.
Our current implementation of Minimmit, however, uses conservative erasure coding parameters. Since Minimmit requires only $2f+1$ messages to progress through views, we could potentially require $2f+1$ valid fragments for view progression and design the erasure coding to allow reconstruction from exactly $2f+1$ fragments rather than the $f+1$ fragments currently required. This would reduce fragment sizes by approximately 50%, yielding lower transaction latency than Kudzu by combining reduced communication overhead with our faster view progression mechanism. We leave coding parameter optimization for future work.
Increasing per processor bandwidth from 1 Gbps to 10 Gbps under this naïve coding scheme, as illustrated in Table 7, restores Minimmit’s outright transaction latency advantage over Kudzu under this workload.
@@
Minimmit $142.14\pm 21.36$ ms $216.27\pm 28.58$ ms $358.41\pm 7.07$ ms
With more bandwidth to compensate for the larger fragment sizes, Minimmit decreases transaction latency by 25.0% relative to Simplex and 9.07% relative to Kudzu. Increasing bandwidth and reducing latency in this standardised environment, as illustrated in Table 7 and Table 4, consistently improves Minimmit’s performance relative to Simplex and Kudzu. This observation supports the primary insight employed by Minimmit (and Simplex): faster-than-finality view progression is a fundamental building block for minimising transaction latency.

### v7

@@ 6.3. Erasure coding
As explained in Appendix C, a naive approach to the integration of erasure codes causes the resulting version of Minimmit to have a data expansion rate of 5: If blocks are of size $B$, this means the total amount of block data sent by each leader (to all processors combined) is approximately $5B$. The factor 5 arises because an M-notarisation only suffices to ensure that $f+1$ correct processors have received their corresponding block fragments, meaning that $f+1$ out of $n$ fragments should suffice for block reconstruction. However, in Appendices C-E, we also describe how to integrate erasure codes with Minimmit to give a data expansion rate of 2.5.
@@ 6.4. Threshold signatures and communication complexity
Section 4 assumes correct processors send newly received transactions to all others. In practice, transactions are typically disseminated through gossip networks, where each processor forwards transactions to a small constant number of peers. If transactions are constant-bounded in size, this approach maintains constant communication overhead per processor per transaction. Alternative dissemination mechanisms like Narwhal (Danezis et al., 2022) can also be employed. In our analysis here, we treat transaction propagation as a black box and focus on the consensus layer. We assume blocks have constant-bounded size and, following standard practice, analyse only the communication costs required for consensus, taking mempool formation as given.
@@
Communication complexity. Each notarisation and nullification contains $\Omega(n)$ signatures, resulting in communication complexity greater than $O(n^{2})$ per view. The standard approach of using threshold signatures (Boneh et al., 2001; Shoup, 2000) can be used to ensure that the communication complexity per view is $O(n^{2})$. We define:
  • An ‘M-certificate’ for $b$ to be threshold signature (of constant-bounded length for a given security parameter) formed from $2f+1$ votes for $b$ by different processors.
  • An ‘N-certificate’ for view $v$, to be a threshold signature formed from $2f+1$ nullify(v)(v) messages by different processors.
Threshold signature implemetation details. If $p_{i}$ has not already formed or received an M-certificate for $b$, then, upon receipt of $2f+1$ votes for $b$ by different processors, $p_{i}$ combines the $2f+1$ signatures to form an M-certificate (a single threshold signature). Rather than storing and forwarding the $2f+1$ votes, $p_{i}$ stores and forwards the M-certificate. Similarly, $p_{i}$ stores and sends N-certificates, rather than storing and sending large collections of nullify(v)(v) messages. Since partial signatures for any processor can be derived from $2f+1$ partial signatures, we must now stipulate that processors finalise a block $b$ upon receiving nfn-f votes for $b$ directly from the corresponding processors.
@@ 7. Experiments
Latency and bandwidth model. For every message transmission we sample a delay from a normal distribution with mean equal to the p50181818https://www.cloudping.co/api/latencies?percentile=p_50&timeframe=1Y AWS inter-region latency and standard deviation $(\text{p90}-\text{p50})$. [^19] https://www.cloudping.co/api/latencies?percentile=p_90&timeframe=1Y Each processor has symmetric 1 Gbps (125,000,000 B/s) ingress and egress budgets, so message delivery time is the sum of the sampled network delay and the bandwidth-limited transmission time. We assume linear bandwidth usage and max-min fairness for bandwidth allocation, yielding conservative latency estimates. In practice, burstable bandwidth and traffic prioritisation provided by cloud operators would only decrease the reported latencies.
@@
Introducing a Reed–Solomon erasure coding scheme reduces this bottleneck. We split each block into 50 fragments, have the leader broadcast a fragment to every processor, and require replicas to broadcast their fragment to all other processors when casting a vote. The fragment size is determined by the quorum needed for view progression: Simplex targets $3f+1$ replicas and therefore transmits 61.72 KB fragments so that any $f+1$ fragments suffice for reconstruction; E-Minimmit (as described in Appendices B-E targets $5f+1$ replicas and sends 55.19 KB fragments to ensure $2f+1$ fragments suffice; similarly, Kudzu also targets $5f+1$ replicas and requires 55.19 KB fragments to guarantee reconstruction from $2f+1$ fragments. Table 6 reports the resulting latencies, increasing Minimmit’s effective processing rate from 10,000 to around 30,000 transactions per second at a lower transaction latency.
@@
Minimmit $177.06\pm 21.45$ ms $251.29\pm 28.63$ ms $428.35\pm 7.08$ ms
With coding enabled, Minimmit achieves the lowest view latency, block latency and transaction latency. Minimmit maximises throughput across all protocols, since blocks of a fixed size are produced at a faster rate.
Increasing per processor bandwidth from 1 Gbps to 10 Gbps under this naïve coding scheme, produces latencies illustrated in Table 7.
@@
Minimmit $138.02\pm 21.37$ ms $212.3\pm 28.6$ ms $350.32\pm 7.07$ ms
Minimmit decreases transaction latency by 26.7% relative to Simplex and 11.1% relative to Kudzu, and also decreases view latency by 26.0% relative to Simplex and 24.1% relative to Kudzu.